# U-index: A Universal Indexing Framework for Matching Long Patterns

### v1

@@ 5 Experiments
Implementation Details. In practice, we verify each candidate occurrence using a linear scan of $P$ in $T$, hence without using any trie data structure. Even if this solution costs $\mathcal{O}(m)$ time instead of the $\mathcal{O}(1)$-time verification claimed by Theorem 4.1, this is likely to be faster in practice because traversing tries is not cache-efficient.
@@ 5.1 Setup
Hardware and Software. All experiments were run on an Intel Core i7-10750H running at a fixed frequency of 2.6 GHz with hyperthreading disabled and cache sizes of 32 KiB (L1), 256 KiB (L2), and 12 MiB (shared L3). Code was compiled using rustc 1.85.0-nightly and GCC 14.2.1 on Arch Linux.
Datasets. We use three textual datasets of different nature and alphabet size: (1) chromosome 1 of CHM13v2.0333https://www.ncbi.nlm.nih.gov/nuccore/NC_060925.1, which contains repetitive regions and consists of 248 million symbols over the DNA alphabet ($\sigma=4$), thus 59 MiB when each symbol is coded using 2 bits; (2) the 200 MiB protein sequences available from the Pizza & Chili site [^4] https://pizzachili.dcc.uchile.cl/texts/protein ($\sigma=27$), or 125 MiB when each symbol is coded using 5 bits; and (3) the 200 MiB English collection, also available from the Pizza & Chili site [^5] https://pizzachili.dcc.uchile.cl/texts/nlang ($\sigma=239$). In the following, we discuss experimental results referring to Figure 4 (at page 4) for the DNA alphabet. The results for proteins and English have very similar shapes and are deferred to Appendix B due to space constraints.
Queries. For each dataset, we test $10^{5}$ positive queries that are uniformly sampled from the text. For DNA, queries consist of $512$ characters, while for the protein and English datasets we use queries of $128$ characters.
Tested Indexes. We compare the suffix array and the FM-index, indicated with SA and FM-index in our results, against their U-index variants, with parameters $(k,\ell)\in\{(4,32),(8,64),(16,128),(28,256)\}$. For the suffix array construction, we use libsais [^6] https://github.com/IlyaGrebnov/libsais [30, 22]. For the FM-index, we use the implementations in SDSL-lite (SDSL-v2) [15] and in AWRY [1].
For each index, we use the smallest $\tau\geq\lceil\log_{2}c\rceil$ that is supported by the index. In practice, this means that for the suffix array and SDSL FM-index we use $\lceil\log_{2}c/8\rceil$ bytes to represent each ID, so that each ID is a single character. This way, these indexes get built on exactly $z=|\mathcal{M}_{\ell,k}(T)|$ characters. In practice, this is strictly better than using a smaller $\tau$ and building an index on $2z$ or more characters. The AWRY FM-index does not support generic alphabets, and thus we had to consistently use $\tau=2$ to encode the IDs as DNA bases. We note that the AWRY uses a multi-threaded parallel construction algorithm, while all other methods are single-threaded. Further details can be found in Appendix A.
@@ 5.2 Results
Figure 4: Results on the $59$ MiB of the human chromosome 1. For each data structure — except the sparse SA — we compare its performance when constructed on the plain input text (in red, left column of each group) versus when used with the U-index (remaining colors and columns), for increasing values of $k$ and $\ell$. Indexes marked with -H (read, “minus H”) use minimizers themselves as IDs, without the map $H$. Similarly, the indexes marked with -S omit storing the sketched input text $S$ and instead reconstruct it via the minimizer positions $\mathcal{M}_{\ell,k}(T)$ and $T$ itself. The sparse SA is only shown with sampling (no red column) because it is otherwise equivalent to SA. The top plot shows space usage of the final data structure in MiB, with the space for minimizer positions and the map $H$ shaded and the black line indicating the space occupied by the 2-bit packed input text. The second row shows the maximum memory usage (resident set size, RSS) during the construction, where the shaded area is memory usage before construction. The third row shows build time (in seconds), with the time for sketching the input shaded. The bottom plot shows query time (in average $\mu$s per Locate query), with the time for searching in the inner index shaded.
@@
Suffix Array. In all cases, we see that the U-index variants take less space and are faster to construct than the classic indexes. Both space and construction time tend to be at least 8×8$\times$8 × less. The time spent searching the suffix array (bottom, shaded) goes down as the suffix array becomes sparser, but the total query time goes up. This is due to the increasing number of false positive matches in minimizer space, starting at $10$ false positives per query for $(k,\ell)=(4,32)$ and going up to $100$ for $(k,\ell)=(28,256)$. These are caused by highly repetitive regions in the centromere [17]. Nevertheless, query time is usually less than twice slower than the classic suffix array, while space and construction time are greatly reduced.
@@
Omitting $H$. When $k\in\{16,28\}$, most minimizers are unique, and the map $H$ mapping the minimizers to IDs has size linear in their number. This can be seen in that the shaded area in the rightmost two columns of the top-left group is accountable for most of the space used by the index at over $642^{6}=$ 64 MiB, while the input only takes 59 MiB when encoded using $2$ bits per character. Looking at SA -H, which omits $H$, we see that the index is much smaller for larger $k$, while query time is unaffected. Construction time does increase, since the alphabet is larger and hence requires more memory. The high spike for $k=16$ seems to be caused by SA (libsais) being particularly slow for 32-bit integers.
@@
Implicit Sketched Text. We can also omit the sketched text $S$ and instead reconstruct it on-the-fly, as explained in Section 4. This saves significant space when $k$ is large (where also $H$ can be omitted altogether). Construction time is not affected since $S$ is discarded after construction. The time spent in searching the suffix array does increase significantly though, up to 4×4$\times$4 × (shaded), but as most of the time is spent verifying potential matches, the total query time only goes up slightly.
@@
FM-index. The SDSL implementation takes around 90 MiB for the plain input text, and goes significantly below this when using the U-index [^7] For some technical reason, SDSL does not accept NULL bytes, hence minimizers must necessarily mapped by $H$ in the range $[1,c+1)$. Thus SDSL takes a significant amount of space for $k\geq 16$.. Construction time improves as well, almost 2×2$\times$2 × each time we double $\ell$ because the number of sampled positions nearly halves.
@@
A main drawback of the SDSL implementation is its significantly larger query times compared to all other methods, starting at 32×32$\times$32 × slower for the plain index and increasing up to 1000×1000$\times$1000 × slower for $\ell=256$. We suspect this is due to the inherent complexity in the wavelet tree data structures used to represent the Burrows-Wheeler transform of the text [6], that has increasingly more levels as $k$ increases and hence the number of bits $\tau$ in each $k$-mer ID grows. This makes SDSL an impractical choice in this scenario or, more in general, for applications where size is not a bottleneck but fast queries are the primary concern.
@@
AWRY uses around twice as much space as SDSL, and has a similar construction time, likely because both are limited by the internal suffix array construction. On the other hand, query times for AWRY are significantly faster than SDSL because AWRY is optimized for the DNA alphabet, and the U-index version with $k=4$ is slightly faster than AWRY on the plain text. However, as $k$ and $\ell$ increase, both SDSL and AWRY slow down by roughly a factor of 2 overall. For AWRY we can also omit $H$ and this almost halves the size when $k$ is large, but negates the previously seen speedup in construction time since the sketched input $S$ is significantly larger.
@@
Sparse SA. Lastly, the sparse SA index takes strictly less space than the U-index variant of the suffix array, since it stores strictly less information: only one permutation of the minimizer positions in the original text. It is also significantly faster to query, since no sketching is needed. Additionally, it has much fewer false positives, since comparisons are made in the plain input space rather than in sketch-space. Construction time is worse than SA however, since there is no known linear-time implementation for sparse suffix array construction [^8] Indeed, our implementation is very simple and relies on sorting of the minimizer positions using the Rust sort_by function that returns the suffix associated with each position as needed.. Nevertheless, this is still faster than constructing an FM-index.
@@ 5.3 An Example Application
Setup. We run the following experiment. As input, we take the full human genome (CHM13v2.0) and a prefix of 450 PacBio HiFi long reads from the HG002-rep1 dataset [^9] Downloaded from https://downloads.pacbcloud.com/public/revio/2022Q4/HG002-rep1/.. These reads are approximately 99.9% accurate and have an average length of 16kbp.
@@
Results. We build the U-index on top of the libsais suffix array in the configuration where the minimizer space sequence $S$ and map $H$ are both stored. We use $k=8$ and $\ell=128$, so that at least 2 and on average 4 minimizers are sampled from every pattern. This results in 53M minimizers, and the entire U-index is built in 12 seconds.
Out of the 450 sequences, 445 have at least one matching pattern, and in total, 14 824 of the 28 243 patterns match (52%). As observed before, an issue with DNA is that it contains many long repetitive regions. In particular, 160 of the patterns each match around 3820 times for a total of 611k matches, while the 28k remaining patterns only match 27k in total. Worse, there are 721M mismatches, i.e., candidate matches in sketch-space that turn out not to be matches in the genome. Verifying these candidate matches takes over 98% of the time.
Limiting Matches. Only very few (<200) patterns have more than 10 matches, and thus we stop searching once we hit 10 matches. Further, most patterns have relatively few mismatches, while a few patterns have a lot of mismatches. To also avoid those negative effects, we generally only consider the first 100 matches in sketch-space. We still match 445 of 450 reads, while the number of matched patterns goes down to 13 717. On the other hand, the number of mismatches is now 663k (23 per query) and the number of matches is 25k (0.9 per query).
@@ 6 Conclusions and Future Work
The sparse suffix array index by Grabowski and Raniszewski [16] remains a great solution in this regime, having smaller size and significantly faster queries than the U-index around suffix arrays (which have, however, faster construction time). For example, albeit somewhat larger than the SDSL FM-index, the sparse suffix array is over 100×100$\times$100 × faster to query. However, the benefit of the U-index framework lies in its universality and usability. We remark that the primary objective of this work is to highlight these two important properties.
We anticipate that the U-index may be especially useful around the r-index [14] when used on highly repetitive data, but we leave this as future work. The sparse suffix array will by design be unable to take advantage of the underlying repetitiveness. Hence, other than universality, another important virtue of our framework is that it preserves string similarity: for any two highly similar texts $T_{1}$ and $T_{2}$, it will be the case that $S_{1}=\textup{{Sketch}}(T_{1})$ and $S_{2}=\textup{{Sketch}}(T_{2})$ are also highly similar (assuming the sketches are based on minimizers).
@@
Appendix A Further Details on Tested Indexes and the $\tau$ Values
Here we explain in more detail how each of the indexes was used.
The libsais Suffix Array. For the plain input, including for DNA, we use a $\tau=8$ byte encoding. For the sketched text $S$, this depends on $c$. When $\log_{2}(c)\leq 8$, we use one-byte encoding and call the libsais function. When $\log_{2}(c)\leq 16$, we use two-byte encoding and call libsais16. For larger $c$, we first remap all IDs to values starting at 00, as recommended by the library authors, and then call the function libsais_int on 32-bit input values.
The SDSL-lite FM-index. For the plain text, we use the Huffman-shaped wavelet tree with default parameters, i.e., the class csa_wt<wt_huff<rrr_vector<63>>,32,64>. For the U-index counterpart, we instead use csa_wt<wt_int<rrr_vector<63>>,32,64>, a wavelet-tree over a variable-width integer alphabet. Both these indexes use a sampling factor of $32$ to sample the suffix array. Also, for both versions we remap $k$-mers to ids starting at $1$ instead of 00, since SDSL does not support input values of 00. For this reason, we do not have a -H no-remap variant of the SDSL FM-index.
The AWRY FM-index. The AWRY FM-index only supports DNA and protein alphabets. For consistency, we only use the DNA version. This means that we consistently use $\tau=2$. Thus, IDs with $\log_{2}(c)$ bits are encoded into $\lceil\log_{2}(c)\rceil$ DNA bases, which are passed to AWRY as 8-bit ACGT characters. Similarly, plain text protein and English input is encoded into 4 underlying characters. This index also uses a suffix array sampling factor of $32$.
The sparse SA. The sparse SA consists of an array of 32-bit integers indicating text positions. It is constructed using the Rust standard library Vec<u32>::sort_unstable_by_key function that compares text indices by comparing the corresponding suffixes.

### v3

@@ 5 Experiments
Implementation Details.
In our implementation, we verify each candidate occurrence using a linear scan of $P$ in $T$, hence without using any trie data structure. Even if this solution costs $\mathcal{O}(m)$ time instead of the $\mathcal{O}(1)$-time verification claimed by Theorem 4.1, this is likely to be faster in practice because (as it is well known) traversing tries is not cache-efficient.
@@ 5.1 Setup
Hardware and Software.
All experiments were run on an Intel Core i7-10750H running at a fixed frequency of 2.6 GHz with hyperthreading disabled and cache sizes of 32 KiB (L1), 256 KiB (L2), and 12 MiB (shared L3). Code was compiled using rustc 1.85.0-nightly and GCC 14.2.1 on Arch Linux.
Datasets.
We use three textual datasets of different nature and alphabet size: (1) chromosome 1 of CHM13v2.0333https://www.ncbi.nlm.nih.gov/nuccore/NC_060925.1, which contains repetitive regions and consists of 248 million characters over the DNA alphabet ($\sigma=4$), thus 59 MiB when each character is coded using 2 bits; (2) the 200 MiB protein sequences available from the Pizza & Chili site [^4] https://pizzachili.dcc.uchile.cl/texts/protein ($\sigma=27$), or 125 MiB when each character is coded using 5 bits; and (3) the 200 MiB English collection, also available from the Pizza & Chili site [^5] https://pizzachili.dcc.uchile.cl/texts/nlang ($\sigma=239$). In the following, we discuss experimental results referring to Figure 4 (on page 4) for the DNA alphabet. The results for proteins and English have very similar shapes and are deferred to Appendix B due to space constraints.
Queries.
For each dataset, we test $10^{5}$ positive queries that are uniformly sampled from the text. For DNA, queries consist of $512$ characters, while for the protein and English datasets, we use queries of $128$ characters.
Tested Indexes.
We compare the suffix array and the FM-index, indicated with SA and FM-index in our results, against their U-index variants, with parameters $(k,\ell)\in\{(4,32),(8,64),(16,128),(28,256)\}$. For the suffix array construction, we use libsais [^6] https://github.com/IlyaGrebnov/libsais [33, 23]. For the FM-index, we use the implementations in SDSL-lite (SDSL-v2) [16] and in AWRY [1].
For each index, we use the smallest $\tau\geq\lceil\log_{2}c\rceil$ that is supported by the index. In practice, this means that for the suffix array and SDSL FM-index we use $\lceil\log_{2}c/8\rceil$ bytes to represent each ID, so that each ID is a single character. This way, these indexes get built on exactly $z=|\mathcal{M}_{\ell,k}(T)|$ characters. In practice, this is strictly better than using a smaller $\tau$ and building an index on $2z$ or more characters. The AWRY FM-index does not support generic alphabets, and thus we had to consistently use $\tau=2$ to encode the IDs as DNA bases. We note that the AWRY FM-index uses a multi-threaded parallel construction algorithm, while all other methods are single-threaded. Further details can be found in Appendix A.
@@ 5.2 Results
Figure 4: Results on the $59$ MiB of the human chromosome 1. For each data structure — except the sparse SA — we compare its performance when constructed on the plain input text (in red, left column of each group) versus when used with the U-index (remaining colors and columns), for increasing values of $k$ and $\ell$. Indexes marked with -H (read, “minus H”) use minimizers themselves as IDs, without the map $H$. Similarly, the indexes marked with -S omit storing the sketched input text $S$ and instead reconstruct it via the minimizer positions $\mathcal{M}_{\ell,k}(T)$ and $T$ itself. The sparse SA is only shown with sampling (no red column) because it is otherwise equivalent to SA. The top plot shows the space usage (size) of the final data structure in MiB, with the space for minimizer positions and the map $H$ shaded, and the black line indicating the space occupied by the 2-bit packed input text. The second row shows the maximum memory usage (resident set size) during the construction, where the shaded area is the memory usage before construction. The third row shows the construction time (in seconds), with the time for sketching the input shaded. The bottom plot shows the query time (in average $\mu$s per Locate query), with the time for searching in the inner index shaded.
Suffix Array.
@@
In all cases, the U-index variants take less space and are faster to construct than the classic indexes. Both space and construction time tend to be at least 8×8$\times$8 × less. The time spent searching the suffix array (bottom, shaded) goes down as the suffix array becomes sparser, but the total query time goes up. This is due to the increasing number of false positive matches in minimizer space, starting at $10$ false positives per query for $(k,\ell)=(4,32)$ and going up to $100$ for $(k,\ell)=(28,256)$. These are caused by highly repetitive regions in the centromere [18]. Nevertheless, the query time is usually less than 2×2$\times$2 × slower than the classic suffix array, while the space and construction time are greatly reduced.
Omitting $H$.
@@
When $k\in\{16,28\}$, most minimizers are unique, and the map $H$ mapping the minimizers to IDs has size linear in their number. This can be seen in that the shaded area in the rightmost two columns of the top-left group is accountable for most of the space used by the index at over $642^{6}=$ 64 MiB, while the input only takes 59 MiB when encoded using $2$ bits per character. Looking at SA -H, which omits $H$, we see that the index is much smaller for larger $k$, while the query time is unaffected. The construction time does increase though, since the alphabet is larger and hence requires more memory. The high spike for $k=16$ seems to be caused by SA (libsais) being particularly slow for 32-bit integers.
Implicit Sketched Text.
@@
We can also omit the sketched text $S$ and instead reconstruct it on-the-fly, as explained in Section 4. This saves significant space when $k$ is large (where also $H$ can be omitted altogether). The construction time is not affected since $S$ is discarded after the construction. The time spent in searching the suffix array does increase significantly though, up to 4×4$\times$4 × (shaded), but as most of the time is spent verifying potential matches, the total query time only goes up slightly.
FM-index.
@@
The SDSL implementation takes around 90 MiB for the plain input text, and goes significantly below this when using the U-index [^7] For some technical reason, SDSL does not accept NULL bytes, hence minimizers must necessarily be mapped by $H$ in the range $[1,c+1)$. Thus, SDSL takes a significant amount of space for $k\geq 16$.. The construction time improves as well, almost 2×2$\times$2 × each time we double $\ell$ because the number of sampled positions nearly halves.
@@
A main drawback of the SDSL implementation is its significantly larger query times compared to all other methods, starting at 32×32$\times$32 × slower for the plain index and increasing up to 1000×1000$\times$1000 × slower for $\ell=256$. We suspect this is due to the inherent complexity in the wavelet tree data structures used to represent the Burrows-Wheeler transform of the text [7], that has increasingly more levels as $k$ increases and hence the number of bits $\tau$ in each $k$-mer ID grows. This makes SDSL an impractical choice in this scenario or, more in general, for applications where the final size is not a bottleneck but fast queries are the primary concern.
@@
AWRY uses around twice as much space as SDSL, and has a similar construction time, likely because both are limited by the internal suffix array construction. On the other hand, query times for AWRY are significantly faster than SDSL because AWRY is optimized for the DNA alphabet, and the U-index version with $k=4$ is slightly faster than AWRY on the plain text. However, as $k$ and $\ell$ increase, both SDSL and AWRY slow down by roughly a factor of 2 overall. For AWRY we can also omit $H$ and this almost halves the size for large $k$, but negates the previously seen speedup in construction time since $S$ is significantly larger.
Sparse SA.
@@
Lastly, the sparse SA index takes strictly less space than the U-index variant of the suffix array, since it stores strictly less information: only one permutation of the minimizer positions in the original text. It is also significantly faster to query, since no sketching is needed. Additionally, it has much fewer false positives, since comparisons are made in the plain input space rather than in sketch space. The construction time is worse than SA, however, since there is no available linear-time implementation for sparse suffix array construction [^8] Indeed, our implementation is very simple and relies on sorting the minimizer positions using the Rust sort_by function that returns the suffix associated with each position as needed.. Nevertheless, this is still faster than constructing an FM-index.
@@ 5.3 An Example Application
Setup.
We run the following experiment. As input, we take the full human genome (CHM13v2.0) and a prefix of 450 PacBio HiFi long reads from the HG002-rep1 dataset [^9] Downloaded from https://downloads.pacbcloud.com/public/revio/2022Q4/HG002-rep1/.. These reads are approximately 99.9% accurate and have an average length of 16kbp.
@@
Results.
We build the U-index on top of the libsais suffix array in the configuration where the sketched text $S$ and map $H$ are both stored. We use $k=8$ and $\ell=128$, so that at least 2 and on average 4 minimizers are sampled from every pattern. This results in 53M minimizers, and the entire U-index is built in 12 seconds.
Out of the 450 reads, 445 have at least one matching pattern, and in total, 14 824 of the 28 243 patterns match (52%). As observed before, an issue with DNA is that it contains many long repetitive regions. In particular, there are 160 patterns that match around 3820 times each for a total of 611k matches, while the 28k remaining patterns only match 27k in total. Worse, there are 721M mismatches, i.e., candidate matches in sketch space that turn out not to be matches in the genome. Verifying these candidate matches takes over 98% of the time.
Limiting Matches.
Only very few (<200) patterns have more than 10 matches, and thus we stop searching once we hit 10 matches. Further, most patterns have relatively few mismatches, while a few patterns have a lot of mismatches. To also avoid those negative effects, we generally only consider the first 100 matches in sketch space. We still match 445 of 450 reads, while the number of matched patterns goes down to 13 717. On the other hand, the number of mismatches is now 663k (23 per query) and the number of matches is 25k (0.9 per query).
@@ 6 Conclusions and Future Work
The sparse suffix array index by Grabowski and Raniszewski [17] remains a great solution in this regime, having a smaller size and significantly faster queries than the U-index around suffix arrays (which have, however, faster construction time). For example, albeit somewhat larger than the SDSL FM-index, the sparse suffix array is over 100×100$\times$100 × faster to query. However, the benefit of the U-index framework lies in its universality and usability. We remark that the primary objective of this work is to highlight these two important properties.
We anticipate that the U-index may be especially useful around the r-index [15] when used on highly repetitive data, but we leave this as future work. The sparse suffix array will by design be unable to take advantage of the underlying repetitiveness. Hence, other than universality, another important virtue of our framework is that it preserves string similarity: for any two highly similar texts $T_{1}$ and $T_{2}$, it will be the case that $S_{1}=\textup{{Sketch}}(T_{1})$ and $S_{2}=\textup{{Sketch}}(T_{2})$ are also highly similar (e.g., assuming the sketches are based on minimizers).
@@
Appendix A Further Details on the Tested Indexes
In this section, we explain in more detail how each of the indexes was used.
The libsais Suffix Array.
For the plain input, including for DNA, we use a $\tau=8$ byte encoding. For the sketched text $S$, this depends on $c$. When $\log_{2}(c)\leq 8$, we use one-byte encoding and call the libsais function. When $\log_{2}(c)\leq 16$, we use two-byte encoding and call libsais16. For larger $c$, we first remap all IDs to values starting at 00, as recommended by the library authors, and then call the function libsais_int on 32-bit input values.
The SDSL-lite FM-index.
For the plain text, we use the Huffman-shaped wavelet tree with default parameters, i.e., the class csa_wt<wt_huff<rrr_vector<63>>,32,64>. For the U-index counterpart, we instead use csa_wt<wt_int<rrr_vector<63>>,32,64>, a wavelet tree over a variable-width integer alphabet. Both these indexes use a sampling factor of $32$ to sample the suffix array. Also, for both versions we remap $k$-mers to IDs starting at $1$ instead of 00, since SDSL does not support input values of 00. For this reason, we do not have a -H no-remap variant of the SDSL FM-index.
The AWRY FM-index.
The AWRY FM-index only supports DNA and protein alphabets. For consistency, we only use the DNA version. This means that we consistently use $\tau=2$. Thus, IDs with $\log_{2}(c)$ bits are encoded into $\lceil\log_{2}(c)\rceil$ DNA characters, which are passed to AWRY as 8-bit ACGT characters. Similarly, plain text protein and English input is encoded into 4 underlying characters. This index also uses a suffix array sampling factor of $32$.
The sparse SA.
The sparse SA consists of an array of 32-bit integers indicating text positions. It is constructed using the Rust standard library Vec<u32>::sort_unstable_by_key function that compares text indices by comparing the corresponding suffixes.